OpenStudy (anonymous):

Calculate the molarity (M) of a solution of the strong acid HBr if the hydroxide ion concentration is 5.00E-12 M.

OpenStudy (aaronq):

For strong acids, we assume that there is 100% ionization, that is: \(\sf HBr\rightarrow H^++Br^-\) We need to find the concentration of \([H^+]\) in the solution, which is equal to \([HBr]\) \(\sf \Large K_{water}^{25^oC}=[H^+]*[OH^-]\) Since you're not told a temperature, we can assume that it's 25 Celsius, and use the value: \(\sf \Large K_{water}^{25^oC}=1.0*10^{-14}~M^2\) We were given \([OH^-]\) in the question: \([H^+]=\dfrac{1.0*10^{-14}}{[OH^-]}=\dfrac{1.0*10^{-14}M^2}{5.00*10^{-12}M }=0.002~M=[HBr]\) The concentration (Molality) of the acid is 0.002 M